# DrivAer Transformer: A high-precision and fast prediction method for vehicle aerodynamic drag coefficient based on the DrivAerNet++ dataset

### v3

@@ I Introduction
With the rapid development of electric vehicles in China, automakers are launching new models faster and faster, and the vehicle development cycle is shortened, which puts forward higher requirements for efficient and accurate aerodynamic calculations. However, traditional automotive aerodynamics relies on the combination of computational fluid dynamics (CFD) and wind tunnel tests, which is a less efficient method and cannot meet the needs of automobile companies to reduce the development cycle and budget. Moreover, CFD requires a high level of experience from development engineers, and in some cases, invalid calculation schemes often occur, and each high-fidelity CFD simulation may take a long time aultman2022evaluation ; biswas2019development ; jiang2021vehicle . As to wind tunnel tests, although the accuracy and reliability is quite good, it can only be used to test a small number of designs due to the need to create an equivalent scale model of the whole vehicle, which increases the time and cost consumed by the experiment. Deep learning-driven approaches can make it possible to speed up the design process and effectively evaluate aerodynamic designs by utilizing existing datasets to explore the aerodynamic performance of different vehicle designs baque2018geodesic ; remelli2020meshsdf ; rios2019scalability ; rios2021point2ffd ; song2023surrogate ; trinh20243d ; benjamin2025systematic .
@@
Owing to intensive research in the field of computer vision, a variety of powerful deep learning techniques designed for processing three-dimensional (3D) models have been developed atzmon2018point ; guo2021pct ; li2018pointcnn ; qi2017pointnet ; qi2017pointnet++ ; wang2019dynamic . These studies have introduced advanced functionalities such as object recognition, classification, and the automatic construction of new models. By leveraging deep learning as an innovative approach, researchers are now able to evaluate aerodynamic performance more efficiently. Through the use of existing model data, they can construct datasets and carry out training, thereby significantly reducing the time and cost typically associated with traditional CFD simulations remelli2020meshsdf ; ribeiro2020deepcfd ; rios2019scalability ; song2023surrogate .
However, current technology has to face a large number of challenging problems in processing three-dimensional (3D) models with higher complexity. For example, the deployment of large deep learning models is limited by the lack of data sets and training resources. At the same time, there is still a significant gap between the prediction accuracy and generalization ability of smaller networks, which makes it difficult to fully meet the actual needs of current production processes jacob2021deep ; tychola2024deep ; vinodkumar2024deep ; wu2024towards .
A significant amount of research has been conducted on basic two-dimensional (2D) models such as airfoils jun2020application ; thuerey2020deep ; wu2020deep , and several mature deep learning methods have been summarized. Among these, DeepCFD, proposed by Ribeiro ribeiro2020deepcfd , utilizes deep learning models like Convolutional Neural Network (CNN) to quickly simulate the velocity and pressure fields of 2D steady-state laminar flow. The deep learning architecture of CFD Net, proposed by Obiols obiols2020cfdnet , leverages extensive model data to predict fluid characteristics rapidly under various boundary conditions and flow field environments, thereby reducing the computation time required by traditional CFD methods. Although these studies can enhance computational speed by several orders of magnitude compared to traditional CFD methods while maintaining lower error rates, they primarily focus on 2D steady laminar flow. There is a scarcity of research related to real-life applications involving complex models. In an effort to advance research, Song proposed proxy modeling of a 3D car model using a novel 2D representation to predict the drag coefficient song2023surrogate . Jacob introduced a different approach by employing a deep learning model to predict the aerodynamic drag coefficient of arbitrary vehicle shapes, addressing a limitation of traditional methods jacob2021deep . This method utilizes an improved U-Net architecture combined with a Signature Distance Field (SDF) to represent the input geometry, enabling accurate predictions of both drag coefficient and velocity fields after training. However, it was noted that the computational cost of augmenting the geometry is high. Given the current computational resources, it may not be practical for the network to learn to recognize the same geometric features from different directions in large-scale datasets. Meanwhile, Charles R. Qi proposed the PointNet architecture, a deep learning network specifically designed for processing 3D point cloud data qi2017pointnet . Its key innovation involves directly handling unordered point sets by processing each point independently and using symmetric functions (e.g., max-pooling) to extract global features, ensuring the network remains insensitive to the arrangement of the input points while remaining invariant to rigid-body transformations. This is significantly different from traditional convolutional neural networks, which rely on structured input data (e.g., images), and enables the extraction of model-specific parameters by analyzing 3D point cloud data. Dynamic Graph Convolutional Neural Network (DGCNN), proposed by Wang wang2019dynamic , is a specialized deep learning framework for 3D point cloud data processing. Its core idea is to characterize the point cloud data by graph structure and introduce dynamic graph convolutional operations to capture local and global geometric features. Unlike traditional static graph methods, DGCNN dynamically updates neighborhood relationships at each layer, allowing the model to flexibly learn complex geometric relationships within point cloud data at multiple scales. The fundamental EdgeConv operation further enhances modeling of local geometric information given the point cloud, effectively improving the accuracy and efficiency of feature extraction. DGCNN has been widely used in classification, segmentation and object recognition tasks involving 3D point cloud data, demonstrating superior performance in handling irregular geometric data. Subsequently, Rios investigated the use of 3D point cloud self-encoders to extract local geometric features in vehicle design optimization rios2021point2ffd . It was found that 3D point cloud self-encoders provide more efficient shape generation and improve the complementary degrees of freedom of the model, leading to improved optimization of aerodynamic performance.
Recently, Elrefaie proposed RegDGCNN, an extended version for regression tasks based on DGCNN elrefaie2024drivaernet . Unlike DGCNN, for the classification task, RegDGCNN employs a regression loss function (e.g., mean square error loss) to optimize the model for scenarios requiring continuous value predictions. This adaptation allows RegDGCNN to excel in applications such as aerodynamic parameter prediction, which demand precise numerical regression. By inheriting the dynamic graph convolution structure of DGCNN, RegDGCNN achieves both efficient feature extraction and accurate regression prediction performance when processing point cloud data with complex geometries, thus supporting a wide range of regression tasks. At its core, RegDGCNN processes point cloud data through EdgeConv. Although it performs well in extracting local geometric features, it is less effective in predicting continuous values. While this method shows strong capabilities in capturing local geometric characteristics, its limitation lies in the inability to effectively model the structural features of the entire shape, which may lead to inaccuracies in aerodynamic calculations. After all, aerodynamic coefficients are influenced not only by local morphology but also by the complex interactions between different components of the overall geometry. Further research by Elrefaie demonstrated that although the RegDGCNN architecture may achieve high prediction accuracy for identical vehicle configurations, its performance significantly declines when applied to models outside the training dataset elrefaie2024drivaernet++ . For entirely new models not present in the dataset, its prediction accuracy falls short of that achieved by CFD simulations, as the EdgeConv-based approach cannot fully leverage the available geometric information, resulting in suboptimal performance for complex shapes. This limitation highlights the urgent need for further advancements in the utilization of geometric data to improve predictive accuracy.
In view of the challenges revealed by the above research, this study proposes a rather novel method that combines the cloud-driven attention mechanism (CDA) with the correlation estimation module (CDE). This integrated approach aims to model the inherent morphological information in the disordered cloud structure in a more effective manner, thereby significantly improving the prediction accuracy and ability of various vehicle designs. This research has resulted in the development of a comprehensive system that can predict the aerodynamic drag coefficient of 3D STL vehicle models. Notably, the system can achieve a prediction accuracy close to that of CFD, while greatly reducing the time and resources generally required for traditional simulations. In addition, after optimization, the system can support the prediction of aerodynamic performance for a variety of vehicle designs, thereby further improving the generality and accuracy of the algorithm.
@@ II DrivAerNet++ Dataset
DrivAerNet++ is the largest and most comprehensive multi-modal 3D dataset available, specifically designed for data-driven aerodynamic analysis in automotive engineering elrefaie2024drivaernet++ . It includes 8,000 industry-standard automotive design models, each sample comes with a detailed 3D model (with up to 500,000 surface meshes) and provides extensive 3D flow fields and precise aerodynamic coefficients. This wealth of information enables researchers and engineers to predict vehicle aerodynamic drag metrics more accurately under various design conditions. With a 333% increase in data volume compared to the previous largest publicly available automotive dataset, DrivAerNet++ is the only open database that includes models of tires and chassis, providing a robust foundation for advanced aerodynamic calculations.
Moreover, DrivAerNet++ offers significant advantages over traditional aerodynamic datasets, such as ShapeNet and other open-source datasets song2023surrogate . Traditional datasets typically include only a limited number of vehicle designs and lack the geometric detail necessary for accurate aerodynamic simulations—especially in the modeling of components like wheels and chassis, which are critical for precise prediction. For instance, vehicle models in ShapeNet are often simplified into single-structure representations, failing to capture the complexity and diversity found in real-world vehicle designs. These simplified models tend to exhibit poor generalization performance when applied to different vehicle types and may even introduce bias.
@@ II.1 CFD Simulation Setup
The study utilizes the DrivAerNet dataset generated through computational fluid dynamics (CFD) simulations implemented in the open-source OpenFOAM® framework elrefaie2024drivaernet ; elrefaie2024drivaernet++ ; greenshields2022openfoam . A pressure-velocity coupled steady-state solver ($simpleFoam$) optimized for incompressible turbulent flow analysis was employed. The computational domain, constructed with 1:1 scale representation of the DrivAer fastback vehicle configuration, features dimensions of $5.0L\times 3.5L\times 2.5L$ (where $L$ denotes vehicle length), incorporating symmetry boundary conditions along the $y$-normal plane to enhance computational efficiency.
@@ II.4 Validation Results
$C_{d}=\frac{F_{d}}{\frac{1}{2}\rho u_{\infty}^{2}A_{\text{ref}}}$ (1)
@@ II.5 Feature Extraction
(a) Raw model (with reduced point cloud size)
(b) Model after contour extraction
Figure 2: DrivAer model along the yz-plane
(a) Raw model (with reduced point cloud size)
(b) Model after contour extraction
Figure 3: DrivAer model along the xz-plane
For each cross-section, the Alpha Shape algorithm is applied to automatically identify edge feature points. A curvature-sensitive and adaptive sampling strategy is then utilized: in regions where the curvature radius falls below a predefined threshold—such as the windshield transition zone and rear wing edges—the sampling density is increased by a factor of three compared to flatter areas. Ultimately, 1,000 feature points are extracted per section to form a sparse point cloud that preserves the aerodynamic characteristics of the vehicle.
This method can control a variety of variables such as the number of cross-sectional layers, sampling density, and curvature threshold in this specific way of parameterization. Throughout the implementation process, it is an important prerequisite to retain the overall vehicle shape, and also to ensure that the point density in key feature areas is higher than that of the traditional Poisson sampling method. The experimental results show that this method can significantly simplify the modeling of irrelevant structural details such as interior parts, radiators, and engines. It is important to note that this simplification process does not affect the prediction accuracy, and ultimately provides a more efficient solution for simulating the drag of the entire vehicle in industrial applications.
@@ III DrivAer Transformer (DAT)
In the developmental research of geometric deep learning in recent years, the application of this method to solve hydrodynamic problems of complex geometries has a good research prospectkashefi2022physics ; pfaff2020learning ; rios2019scalability ; rios2021point2ffd ; rios2019efficiency ; sanchez2020learning . In this study, a series of deep learning modules for processing 3D point cloud data are designed and implemented aiming to efficiently predict the aerodynamic drag coefficient $C_{d}$ of a car through techniques such as geometric feature extraction, dynamic convolution and self-attention mechanism. The model is capable of extracting complex geometrical structures and topological information from 3D point clouds, which greatly enhances the accuracy of aerodynamic parameter prediction. The detailed architecture is illustrated in Fig. 4 and the original diagram is showed in appendix.
@@ III.1 A regression prediction framework based on dynamic graph networks
Aerodynamic drag is strongly influenced by the morphology of the vehicle, so the prediction of aerodynamic drag needs to capture the local correlation of the point cloud data and the complex correlation effect between the global point cloud features. However, 3D point cloud data is a kind of disordered 3D data, in order to maintain the consistency of the learning results while using convolutional neural network for feature extraction of disordered point sets, the maximum pooling is used to aggregate the disordered point set features to form the global point set features. Moreover, for the local vehicle structure, the local geometric structure is utilized by applying a convolution-like operation on the edges connecting neighboring point pairs by constructing bowtie graph values, which enables the model to construct the local point cloud into a graph neural network thus capturing the key aerodynamic parameters of the fluid around the object more accurately. In this paper, on the basis of RegDGCNN, we add the point cloud structure correlation-driven attention (CDA) and correlation estimation module (CDE) to further model the morphological information of the disordered point cloud structure, which more accurately exploits the graph neural network features of the local tie points. Meanwhile, the spatial coding capability of PointNet extracted from the original point cloud features and the local relational inference of DGCNN are utilized to achieve the prediction of continuous values of aerodynamic parameters, and the feature bridges extracted by correlation attention perform more prominently in terms of translational invariance and nonlocality . After multiple layers of correlation edge convolution blocks, the constructed local bowtie graph changes in order to dynamically adapt to the changing point set feature space. In this study, the initialization graph $G$ and a point cloud feature set $X=\{x_{1},\ldots,x_{n}\}\in\mathbb{R}^{F}$ are defined. The graph feature $\gamma\in\mathbb{R}^{C\times P}$, where $C$ represents the number of channel dimensions, and $P$ represents the number of points.
@@ III.2 Correlation-Driven Attention(CDA)
Since edge convolution (EdgeConv) can effectively model the geometric features of locally disordered point cloud structures when processing point cloud data, it is especially good at capturing semantic information in local regions, however, it still has some limitations in extracting global semantic structure features, especially for the application of predicting drag coefficient in aerodynamic tasks. This limitation is mainly reflected in the fact that edge convolution mainly focuses on the relationship between local feature points and ignores the global semantic structure of the overall point cloud, while the prediction of air drag coefficient often relies on the overall geometrical shape and the interactions between the parts. Therefore, the dynamic graph model constructed based on edge convolution is limited in its ability to extract global information and cannot fully utilize the global structural features of the point cloud, thus affecting the accurate prediction of aerodynamic performance.
To address this problem, this paper proposes an innovative attention mechanism based on point cloud feature invariance, which aims to further enhance the extraction of global semantic information through matrix transposition operations. In this method, we introduce a learned attention score at the level of channel features to dynamically regulate the importance of features in different channels. Specifically, by weighting the important channels, local features are made to be highlighted or suppressed in specific channels, which can capture the key parts of the global structure of the point cloud more effectively. In this way, the model can not only continue to play the advantage of edge convolution in local feature extraction, but also make up for its shortcomings in global information extraction through the channel attention mechanism.
In addition, the correlation-based attention mechanism proposed in this paper calculates the similarity between local graph features and using the similarity matrix formed between the transposed graph and the original input graph, the model’s ability to capture global geometric semantic information is significantly enhanced. The specific process is to calculate the similarity between the transposed graph features and the input features to generate a similarity matrix, which in turn effectively represents the global relationship between points. In order to further improve the expressive ability of the model, a new activation method is designed in this paper to enhance the expressive ability of the model by adjusting the nonlinear characteristics through the learnable parameters $p_{1}$, $p_{2}$, and $\beta$. This flexibility allows the model to better adapt to different input features. The following formulas illustrate the approach, assuming the AconC operator input $x\in\mathbb{R}^{1\times C\times 1}$ and learnable parameter $p_{1}\in\mathbb{R}^{1\times C\times 1}$:
@@
where $S$ represents the similarity matrix, $\tau$ denotes batch normalization, $\sigma$ is the Sigmoid activation function, and $\gamma^{T}$ represents the transposed input point cloud features.

### v5

@@ I Introduction
With the rapid development of China’s electric vehicle industry, OEMs are launching new models at an accelerated rate, and the vehicle development cycle has been shortened dramatically, placing higher demands on efficient and accurate aerodynamic calculations. However, traditional vehicle aerodynamic evaluation relies on a combination of computational fluid dynamics (CFD) and wind tunnel tests, which is a less efficient method that is difficult to meet the needs of vehicle manufacturers to shorten development cycles and control budgets. In addition, CFD requires a high level of experience from engineers, is prone to invalid calculation schemes, and takes a long time for each high-fidelity simulation aultman2022evaluation ; biswas2019development ; jiang2021vehicle . Although the wind tunnel test has high accuracy, it is difficult to cover a large number of design options due to the need to produce a solid scaled model, which is costly and time consuming. In recent years, deep learning-driven methods have been able to rapidly evaluate the aerodynamic performance of different models by utilizing existing datasets, thus significantly speeding up the design process baque2018geodesic ; remelli2020meshsdf ; rios2019scalability ; rios2021point2ffd ; song2023surrogate ; trinh20243d ; benjamin2025systematic .
@@
Driven by research in the field of computer vision, multiple deep learning methods for processing 3D models have emerged atzmon2018point ; guo2021pct ; li2018pointcnn ; qi2017pointnet ; qi2017pointnet++ ; wang2019dynamic . The ability of object recognition, classification and automatic modeling makes it possible to achieve efficient pneumatic evaluation through deep learning.This type of method can significantly reduce the time and cost required for traditional CFD by building datasets for training with existing model data remelli2020meshsdf ; ribeiro2020deepcfd ; rios2019scalability ; song2023surrogate .
However, existing methods still face many challenges when dealing with complex 3D models. For example, large models are limited by training resources and dataset size, while small models have difficulty in balancing accuracy and generalization ability, limiting practical applicationsjacob2021deep ; tychola2024deep ; vinodkumar2024deep ; wu2024towards .
Current research mostly focuses on simplified models such as 2D airfoils jun2020application ; thuerey2020deep ; wu2020deep . For example, Ribeiro’s DeepCFD ribeiro2020deepcfd uses convolutional neural network (CNN) to quickly simulate 2D laminar velocity and pressure fields. Obiols’ CFD Net obiols2020cfdnet predicts the flow field characteristics under different boundary conditions with a large amount of model data, which effectively reduces the computational overhead. Although these methods have achieved orders of magnitude speedups in 2D steady state problems, they are still limited in handling complex real-world 3D scenarios. Song song2023surrogate predicts the aerodynamic drag coefficients of 3D vehicles with a 2D agent model, and Jacob jacob2021deep combines the improved U-Net and the Signed Distance Field (SDF) to predict the aerodynamic properties of 3D vehicle body, which can effectively improve the accuracy, but the geometric enhancement overhead is large and difficult to be used for large-scale data training.
To deal with 3D point clouds, Qi proposes the PointNet architecture qi2017pointnet , which can directly deal with unordered point sets, extract global features, and possess rigid-body transformation invariance. Wang proposes DGCNN wang2019dynamic to further capture local and global geometric features through dynamic graph structures, and its core EdgeConv operation performs well in classification, segmentation, etc. A follow-up study by Rios has shown that the self-encoder structure can be used to efficiently extract local geometric features in automotive design optimization rios2021point2ffd , which improves shape generation and aerodynamic performance optimization.
Elrefaie’s recently proposed RegDGCNNelrefaie2024drivaernet++ improves the DGCNN for regression tasks by introducing a mean-square error loss function for continuous-value prediction, which is suitable for tasks such as aerodynamic parameter regression. Although RegDGCNN performs well in local feature extraction, it still has deficiencies in modeling the overall structural features, especially when facing the new car models that are not seen in the training set, the generalization ability decreases significantly elrefaie2024drivaernet++ , and the prediction accuracy is not as good as that of CFDs. This exposes the shortcomings of the current model in dealing with the geometry of the complex shapes, which does not make sufficient use of the geometric information. Information is not fully utilized, and new methods are urgently needed to improve it.
To address the above problems, this paper proposes an innovative approach: combining the point cloud attention mechanism (CDA) with a correlation estimation module (CDE), aiming to model the intrinsic geometric information in disordered point cloud structures more effectively. The method implements a complete system capable of accurately predicting aerodynamic drag coefficients based on a 3D STL vehicle model with an accuracy close to that of CFD, while significantly reducing computational resource consumption. After optimization, the system can be adapted to a wide range of vehicle models, which further improves the generalization ability and accuracy of the algorithm.
@@ II DrivAerNet++ Dataset
DrivAerNet++ is the most comprehensive large-scale multimodal dataset for data-driven automotive aerodynamic analysis elrefaie2024drivaernet++ . It contains 8,000 high-fidelity, industry-standard vehicle models, each with up to 500,000 mesh faces, full 3D flow fields, and precise aerodynamic coefficients. With a 333% increase in data volume over previous datasets, it is the only open-source database that includes detailed wheel and chassis geometries, making it ideal for advanced aerodynamic modeling.
Compared to traditional datasets like ShapeNet song2023surrogate , DrivAerNet++ offers significantly higher geometric detail and design diversity. Conventional datasets often use overly simplified vehicle shapes, lacking critical components such as wheels and underbodies, which results in poor simulation accuracy and limited generalization.
@@ II.1 CFD Simulation Setup
The study utilizes the DrivAerNet dataset generated through computational fluid dynamics (CFD) simulations implemented in the open-source OpenFOAM® framework elrefaie2024drivaernet++ ; greenshields2022openfoam . A pressure-velocity coupled steady-state solver ($simpleFoam$) optimized for incompressible turbulent flow analysis was employed. The computational domain, constructed with 1:1 scale representation of the DrivAer fastback vehicle configuration, features dimensions of $5.0L\times 3.5L\times 2.5L$ (where $L$ denotes vehicle length), incorporating symmetry boundary conditions along the $y$-normal plane to enhance computational efficiency.
@@ II.2 Turbulence Modeling
(a) Raw model (with reduced point cloud size)
(b) Model after contour extraction
Figure 2: DrivAer model along the yz-plane
(a) Raw model (with reduced point cloud size)
(b) Model after contour extraction
Figure 3: DrivAer model along the xz-plane
@@ II.4 Validation Results
$C_{d}=\frac{2F_{d}}{\rho u_{\infty}^{2}A_{\text{ref}}}$ (1)
@@ II.5 Feature Extraction
Edge feature points are automatically identified per cross-section using the Alpha Shape method. An adaptive sampling strategy, sensitive to curvature, increases point density threefold in areas with small curvature radii—such as windshield transitions and rear edges—compared to flatter regions. A total of 1,000 feature points per section form a sparse point cloud preserving key aerodynamic characteristics.
This approach provides flexibility in controlling slices, sampling density, and curvature thresholds. It efficiently retains critical aerodynamic geometry, discards irrelevant internal details, and significantly enhances simulation efficiency without compromising prediction accuracy.
@@ III DrivAer Transformer (DAT)
Recent advancements in geometric deep learning have shown promising potential in solving hydrodynamic problems involving complex geometries kashefi2022physics ; pfaff2020learning ; rios2019scalability ; rios2021point2ffd ; rios2019efficiency ; sanchez2020learning . In this study, we design a deep learning framework to predict the aerodynamic drag coefficient ($C_{d}$) of vehicles from 3D point cloud data. The model leverages geometric feature extraction, dynamic graph convolution, and a self-attention mechanism to effectively capture both local and global aerodynamic features. The overall architecture is shown in Fig. 4, with additional visualizations provided in the appendix.
@@ III.1 A regression prediction framework based on dynamic graph networks
Aerodynamic drag is highly sensitive to vehicle geometry, necessitating a model that captures both local correlations and global structural patterns in 3D point cloud data. Since point clouds are unordered, conventional convolutional networks aggregate global features using max pooling to maintain permutation invariance. Locally, a graph-based representation enables convolution-like operations across neighboring points, allowing the model to extract flow-related features more precisely.
Building upon the RegDGCNN architecture, this research introduce a Correlation-Driven Attention (CDA) mechanism and a Correlation-Driven Estimator (CDE) to improve structural feature extraction from disordered point clouds. The hybrid framework combines PointNet’s spatial encoding with DGCNN’s local relationship modeling, enabling continuous regression of aerodynamic coefficients. The attention-guided feature bridge enhances both translation invariance and global expressiveness. Multi-layer correlation edge convolutions adaptively update the local graph structure based on dynamic changes in the input features. Given a point cloud $X=\{x_{1},\ldots,x_{n}\}\in\mathbb{R}^{F}$ and an initial graph $G$, the extracted graph features are denoted $\gamma\in\mathbb{R}^{C\times P}$, where $C$ is the number of channels and $P$ the number of points.
@@ III.2 Correlation-Driven Attention(CDA)
While EdgeConv effectively models local geometric features, it lacks the ability to capture global semantic structures necessary for accurate drag prediction. To address this, we propose a global attention mechanism that leverages point cloud feature invariance. This mechanism performs channel-wise dynamic reweighting to highlight critical structural features using learned attention weights, while preserving EdgeConv’s local modeling advantages.
Specifically, the similarity between transposed graph features and the original input features is computed to generate a similarity matrix, thereby enriching the model’s understanding of global spatial relationships. This formulation enables the network to emphasize globally important regions while maintaining local detail.
To further enhance expressiveness, a novel activation function parameterized (AconC ma2021activate ) by learnable variables $p_{1}$, $p_{2}$, and $\beta$ is introduced, allowing the model to adapt nonlinearity based on input characteristics. For input $x\in\mathbb{R}^{1\times C\times 1}$, the activation operates as follows:
@@
where $S$ represents the similarity matrix, $\tau$ denotes batch normalization, $\sigma$ is the Sigmoid activation function, and $\gamma^{T}$ represents the transposed input point cloud features. The AconC activation function introduces a switching factor to learn the parameter transition between non-linear (activation) and linear (non-activation) states. This allows the model to more flexibly select the appropriate mapping approach based on varying point cloud features.